VEGAS Team Meeting

GB-137-hub
192.33.117.12##7144
434-817-6443# (telephone)

Attendees:

Adam, Marty, Richard, Joe, Paul, Randy, Mark, Ray, Anish, Ray

Agenda:

  1. Any updates on the updated OG(P) calibration
  2. Items tested last week
  3. Test Time
  4. Other

Meeting notes:

  1. Any updates on the updated OG(P) calibration
    • new code for calibration scheme from Dave McMahon?
      • ongoing contact but work is not yet complete
  2. Items tested last week
    1. Integration time MR
      • Not yet ready to release yet
      • HPC server bug requires blanking fix, but needs a real data test to verify
      • Some Mode 10-19 fix dependency
      • Independent of hardware fix
    2. Mode 10-19 hardware fix
      • Fixed?
    3. Baseline ringing (high frequency noise) in VEGAS data
      • Set for now until new cables may be fabricated
    4. VEGAS Blanking fix and impacts
      • Verify new fix still works
  3. Test Time
    • Wednesday
      • 0800 - 1000 Balancing [Mello, Roshi]
      • 1000-1200 [Roshi]
    • Thursday (all day)
      • 1200-1630 Test the LBW modes [Brandt, Creager]
  4. Other

-- MartyBloss - 2014-10-20
Topic revision: r1 - 2014-10-20, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback